7. In a disk with radius R = 30cm that is in equilibrium, two pairs of forces act, F1 = F2 = 10N and F3 = F4 = 15N. Find the arm of the pair of forces F3 F4


1
Expert's answer
2021-03-14T19:15:56-0400
"(F_1+F_2)(R)=(F_3+F_4)(r)\\(20)(30)=(30)(r)\\r=20\ cm"